OpenStudy (genralduck):

simplify the quotient. be sure to rationalize the denominator. do not change a decimal, give an exact answer. please show your work.

OpenStudy (genralduck):

sqrt (12) over 2-sqrt(3)

OpenStudy (mecharv):

Type it using equations!

OpenStudy (mecharv):

is it \[\sqrt{12}\times2\sqrt{3}\]

OpenStudy (mecharv):

is that a minus?

OpenStudy (genralduck):

its a fraction

OpenStudy (genralduck):

so sqrt ()12 is on top

OpenStudy (mecharv):

okay..

OpenStudy (genralduck):

and 2-sqrt(3) is on the bottm

OpenStudy (mecharv):

okay1

OpenStudy (genralduck):

do you know how i could do a problem like this?

OpenStudy (mecharv):

Yes

OpenStudy (mecharv):

are you familiar with rationalization?

OpenStudy (genralduck):

no, its a new concept to me

OpenStudy (mecharv):

Okay so what you have to do is...multiply the numerator and denominator with the conjugate of the denominator.. I will attach a pic wait.

OpenStudy (mecharv):

you have options?

OpenStudy (genralduck):

nope, its a show your work problem

OpenStudy (mecharv):

Okay I am done... I will attach now..

OpenStudy (genralduck):

thanks, that helps alot

OpenStudy (mecharv):

You can also write it as \[2\sqrt{3}\left( 2 +\sqrt{3}\right)\]

OpenStudy (genralduck):

ok, thanks

OpenStudy (genralduck):

thanks for the help